ISOMERIZATION OF CH_{2}Cl-I TO CH_{2}ClI IN CRYOGENIC MATRICES: A STUDY ON ULTRAFAST TIMESCALE

THOMAS J. PRESTON, \underline{MAITREYA DUTTA}, BRIAN J. ESSELMAN, MICHAEL A. SHALOSKI, ROBERT J. MCMAHON and F. FLEMING CRIM, The University of Wisconsin-Madison Department of Chemistry, 1101 University Avenue, Madison, WI, 53706; AMIABLE KALUME, LISA GEORGE and SCOTT A. REID, Department of Chemistry, Marquette University, Milwaukee, WI, 53233.

We follow up on the previous talk on ultrafast timescale studies of the isomerization between CH_{2}ClI and CH_{2}Cl-I in cryogenic matrices. We establish a population of CH_{2}Cl-I in cryogenic matrices and then pump the two lowest electronic absorption features of CH_{2}Cl-I near 435 nm and 800 nm. Then we study the formation of CH_{2}ClI and CH_{2}Cl-I by probing electronic absorption features of both isomers.